OpenStudy (anonymous):

what is the derivative of cos^2x^3

OpenStudy (anonymous):

\[=−6x^2\cos(x^3)\sin(x^3)\]

OpenStudy (anonymous):

Do you mean \[\cos ^{2}(x^3)\]? If you do, then you need to look up the trig identity for derivative of cos^2 and then apply chain rule (if you don't know it you can ask, but it's also googleable.)